# Switching the Loss Reduces the Cost in Offline Reinforcement Learning

### v2

@@ 3 FQI-log: Fitted Q-Iteration with log-loss
Input: Dataset $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i\in[n]}$, function class $\mathcal{F}$, and $k\in\mathbb{N}$.
@@
    $f_{j}\leftarrow\mathop{\rm arg\,min}\limits_{f\in\mathcal{F}}\sum_{i=1}^{n}\ell_{\log}\left(f(S_{i},A_{i});\,C_{i}+\gamma f_{j-1}^{\wedge}(S_{i}^{\prime})\right)$// $\ell_{\log}$ defined in Eq. 2
@@ 4 Theoretical Results
In this section, we present our main theoretical contribution, the first small-cost bound for an efficient algorithm in batch RL. We also state the assumptions under which our analysis holds and provide a sketch for proving the main result. We start by introducing the definition of an admissible distribution.
Definition 4.1 (Admissible distribution).
4.1 Assumptions
We introduce the necessary assumptions under which our main theorem holds.
Assumption 4.2 (Data distribution).
We have that $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ is a sequence of independent, identically distributed random variables such that $(S_{i},A_{i})\sim\mu$, $C_{i}=c(S_{i},A_{i})$ and $S_{i}^{\prime}\sim P(S_{i},A_{i})$.
Assumption 4.3 (Concentrability coefficient).
There exists $C<\infty$ such that any admissible distribution $\nu$ satisfies
$\frac{\nu(s,a)}{\mu(s,a)}\leq C$ (3)
for all $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Assumption 4.4 (Realizability).
$q^{\star}\in\mathcal{F}$.
size=,color=green!20!white,inline]AA: According to Chen and Jiang (2019), for finite $\mathcal{F}$ completeness implies realizability, so this assumption might be redundant
Assumption 4.5 (Completeness).
For all $f\in\mathcal{F}$ we have that $\mathcal{T}f\in\mathcal{F}$.
Assumption 4.6 (Bounded targets).
There exists $B>0$ such that, for all $f\in\mathcal{F}$ and $i\in[n]$, $(\mathcal{T}f)(S_{i},A_{i})$ and $C_{i}+\gamma f^{\wedge}(S_{i}^{\prime})$ are in $[0,B]$ almost surely. size=,color=orange!20!white,]SR: This was a little tricky to read I think, $\mathcal{T}f$ is $\mu$-almost surely but it’s not super obvious that’s what’s meant. I changed to $(\mathcal{T}f)(S_{i},A_{i})$. Also I switched the quantification around, there exists $B$ such that for all … instead of for all … there exists $B$. I’m pretty sure the first version wasn’t what we need.
Assumptions 4.2, 4.3, 4.4 and 4.5 are commonly made, in various forms, when analyzing fitted Q-iteration (Farahmand, 2011; Pires and Szepesvári, 2012; Chen and Jiang, 2019). Assumption 4.3 ensures that all admissible distributions are covered by the exploratory distribution $\mu$ and that $C$ is manageably large. Stated simply, concentrability allows us to assume that $\mu$ sufficiently explores $M$. Assumption 4.4 is the standard realizability assumption which guarantees that the optimal action-value function, our ultimate target, lies in our function class. Assumption 4.5 intuitively states that the function class $\mathcal{F}$ is closed under the Bellman optimality operator $\mathcal{T}$. This assumption is necessary, as a result from Foster et al. (2021) states that Assumptions 4.3 and 4.4 are not sufficient for sample efficient batch value function approximation. Finally, Chen and Jiang (2019) mention that when $\mathcal{F}$ is finite, completeness implies realizability. size=,color=orange!20!white,]SR: It seems very obvious, do we need to cite this fact? Also should we include this since we aren’t really assuming finite $\mathcal{F}$ anymore? For a more detailed discussion of Assumptions 4.3, 4.4 and 4.5, we refer the reader to Sections 4 and 5 of Chen and Jiang (2019).
Assumption 4.6 is novel to this work and thus we discuss it separately. For our analysis, we require both inputs to log-loss, predictor and target, to be in a bounded range. Thus, we require an assumption that the function class $\mathcal{F}$ and MDP $M$ are such that for all $i\in[n]$ the targets $C_{i}+\gamma f^{\wedge}(S_{i}^{\prime})$ lie in a bounded interval. Without loss of generality we let $B=1$ since any bounded domain can be normalized to $[0,1]$.
With our assumptions stated we are ready to present the main theoretical contribution of this work, a small-cost error bound for FQI-log.
4.2 Main Theorem
@@
Given a dataset $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ with $n\in\mathbb{N}$ and a finite function class $\mathcal{F}\subseteq(\mathcal{S}\times\mathcal{A}\rightarrow[0,1])$ that satisfy Assumptions 4.2, 4.3, 4.5 and 4.6, with $B=1$, it holds with probability $1-\delta$ that the output policy of FQI-log after $k$ iterations, $\pi_{k}=\pi_{f_{k}}$, satisfies
$\bar{v}^{\pi_{k}}-\bar{v}^{\star}\lesssim\frac{1}{(1-\gamma)^{2}}\left(\sqrt{\frac{\bar{v}^{\star}C}{n}}+\frac{C}{(1-\gamma)^{2}n}+\gamma^{k}\right)\,.$
The full statement of Theorem 4.7, including lower order terms, can be found in Appendix B along with its proof. Compared to prior error bounds for FQI (Antos et al., 2007, 2008; Munos and Szepesvári, 2008; Farahmand, 2011; Lazaric et al., 2012; Chen and Jiang, 2019), to the best of our knowledge, Theorem 4.7 is the first that contains the instance-dependent optimal cost $\bar{v}^{\star}$. This makes Theorem 4.7 a small-cost bound, also referred to as a first-order (Freund and Schapire, 1997; Neu, 2015) or small-loss (Lykouris et al., 2022; Wang et al., 2023) bound in the learning theory literature. All previous results for FQI obtain an error bound independent of $\bar{v}^{\star}$, and cannot be made to scale with $\bar{v}^{\star}$ due to their use of squared loss. This is the consequence of a somewhat surprising result, Theorem 2 of Foster and Krishnamurthy (2021), which states that a value function learned with squared loss fails to achieve small-cost bounds for cost sensitive classification. Finally, we highlight that Theorem 4.7 is the first small-cost bound for a computationally efficient batch RL algorithm.
Remark 4.8.
We assume the function class $\mathcal{F}$ is finite to simplify our analysis and presentation. For infinite function classes that admit a finite lower bracketing number, our approach can be extending by replacing $\mathcal{F}$ with its lower bracketing cover (Zhang, 2023).
@@ 4.3 Proof Sketch
The proof of Theorem 4.7 is presented in its entirety in Appendix B. We start by defining the pointwise triangular deviation of $f$ from $q^{\star}$,
@@
which is closely related to triangular discrimination (Topsoe, 2000). We can relate $\Delta^{2}_{f}$ to the Hellinger distance via the following lemma.
@@
where for $p=q=0$ we define the left-hand side to be zero and $\operatorname{h}^{2}(p\mathrel{\|}q)=(\sqrt{p}+\sqrt{q})^{2}+(\sqrt{1-p}+\sqrt{1-q})^{2}$ is the squared Hellinger distance.
The proof of Lemma 4.9 is deferred to Lemma A.1. The idea to relate the pointwise triangular deviation to the squared Hellinger distance was first employed by Foster and Krishnamurthy (2021) in analyzing regret bounds for contextual bandits. Our proof can be summarized by the following three main steps, which correspond to the three terms given in Lemma 4.9.
@@
The first step in the proof is to decompose the error (or suboptimality gap), $\bar{v}^{\pi_{k}}-\bar{v}^{\star}$, into the product of a small-cost term and the pointwise triangular deviation of $f$ from $q^{\star}$. The analysis in this step is inspired by the proof of Lemma 1 of Foster and Krishnamurthy (2021). We deviate from their analysis to avoid introducing an extra $|\mathcal{A}|$ factor in the bound. We use the performance difference lemma (Lemma B.4), a multiplicative Cauchy-Schwarz (Lemma B.5), i.e. for distribution $\nu$
@@
and an implicit inequality (i.e. Lemma B.7, step $\star$ in the proof of Proposition B.2) in order to get a small-cost decomposition of the error.
@@
$\bar{v}^{\pi}-\bar{v}^{\star}\lesssim\frac{D_{f}}{1-\gamma}\sqrt{\bar{v}^{\star}}+\frac{D_{f}^{2}}{(1-\gamma)^{2}}\,.$
@@
where $\nu\in\mathcal{M}_{1}(\mathcal{S}\times\mathcal{A})$. Lemma B.13 establishes the contraction properties of $\mathcal{T}$ with respect to the Hellinger distance, i.e.
$\left\|\sqrt{\mathcal{T}f}-\sqrt{\mathcal{T}q^{\star}}\right\|_{2,\nu}\leq\gamma^{1/2}\left\|\sqrt{f}-\sqrt{q^{\star}}\right\|_{2,\nu^{\prime}}$
@@
Contraction arguments have a long history (Bertsekas, 1995; Littman, 1996; Antos et al., 2007; Chen and Jiang, 2019) in the analysis of dynamic programming algorithms that solve MDPs. The novelty is that we need to bound the pointwise triangular deviation. This requires new proofs. Note that such change of measure arguments used in the “error propagation analysis” of RL algorithms are standard.
In Lemma B.15, we use the contraction property to bound the approximation error by
@@
where $\mu$ is the exploratory distribution in Assumption 4.2. In batch RL, we often want to learn $q^{\star}$, or a policy whose value is close to $q^{\star}$. However, when using function approximation FQI, the regressor corresponding to the training data is $\mathcal{T}f$. Therefore, as demonstrated in Eq. 5, if we can show the contraction property then we can bound the error between $f$ and $q^{\star}$ by the error between $f$ and $\mathcal{T}f$. We will argue shortly, the error between $f$ and $\mathcal{T}f$ goes to zero as the size of the batch dataset grows.
Step 3: Error propagation
@@
Then by application of Theorem A.3, we have that if $f$ is the minimizer of $\ell_{\log}$ with respect to the batch dataset $D$, then
$\sqrt{2}\left\|\operatorname{h}^{2}(f\mathrel{\|}\mathcal{T}f)\right\|_{1,\mu}^{1/2}\lesssim\frac{2}{\sqrt{n}}\,.$
We then use the above result and to control the pointwise triangular deviation in Proposition 4.10, i.e $D_{f}$,
@@
This provides a sketch for proving Theorem 4.7. In the full proof we also need the control the individual iterates of FQI-log, i.e. the $f_{j}$’s. We fill in the missing details in our appendix.
@@ 5 Numerical Experiments
The goal of our experiments is to provide insights into the benefits of using FQI-log for learning a near-optimal policy in batch reinforcement learning. We run our first set of experiments in reinforcement learning with linear function approximation, as this setting allows us to best compare FQI-log to FQI-sq without other confounding factors. We then run experiments in deep reinforcement learning to verify whether our findings generalize to more complex function classes. The objective is either to minimize squared loss, which does not get small-cost bounds, or to minimize log-loss, as proposed in this work.
5.1 Measurements
In the first set of experiments, we report the value of the learned policy both as a function of the number of training samples and the number of “good” trajectories in the batch dataset to indicate the sample efficiency of the two algorithms. The results for mountain car are obtained from $90$ independently collected batch datasets, generated by following the uniform random policy. In the second set of experiments, we report the cumulative undiscounted reward, as is standard in deep reinforcement learning (Mnih et al., 2015), on Asterix and Seaquest as a function of the number epochs. The batch datasets for both Asterix and Seaquest are obtained from the $5$ independently collected batch datasets given by Agarwal et al. (2020).
5.2 Mountain Car
@@
Figure 1: The value of the policy learned by FQI as a function of the size of batch dataset. The result are averaged over $90$ independently collected datasets. The figures on the left, middle and right are generated using batch data that contain only $1$, $5$, and $30$ successful trajectories respectively. The standard error of the mean is report via the shaded region.
We first evaluate FQI-log and FQI-sq on an episodic sparse cost variant of Mountain Car (Moore, 1990), where the value functions are indexed by the timestep, i.e. $v_{h}$. This environment consists of a $2$-dimensional continuous state space and $3$ discrete actions. The agent, i.e. the policy $\pi_{f_{k}}$, starts from a fixed state, $s_{1}=[-0.5,0]$, and interacts with the environment for 800 timesteps before an episode terminates. If an agent reaches the top of the hill, it remains there until the end of the episode. The cost is 0 at all timesteps except the last, when a cost of 1 is received if the agent has not reached the hilltop, and otherwise the cost is 0. A discount factor of $\gamma=1$ is used in our experiments.
We encode states as feature vectors given by the Fourier basis of order $2$ (Konidaris et al., 2011), $\phi(s)\in\mathbb{R}^{9}$, as these are known to be good features for value estimation in Mountain Car, see Chapter 9 of Sutton and Barto (2018). Given these features our function class $\mathcal{F}$ for this environment is
$\left\{f_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]:\exists\theta_{a}^{h}\ \text{ s.t. }\ f_{h}(s,a)=\sigma\left\langle\phi(s),\,\theta_{a}^{h}\right\rangle\right\},$
@@
where $\sigma(x)=(1+\exp(-x))^{-1}$ is the sigmoid function and $h\in[800]$. We employ the BFGS method, a quasi-Newton method with no learning rate, to find the minimizer of the losses. For strongly-convex functions, BFGS is known to converge to the global minimum superlinearly (Dennis and Moré, 1974). Finally, each batch dataset is constructed from a set of trajectories collected by running the uniform random policy from the initial state $30,000$ times. We use rejection sampling in order to guarantee each dataset has $[1,5,30]$ trajectories that reach the top of the hill. We train both FQI-log and FQI-sq on the same batch datasets with the first $n=[1,3,6,9,12,15,18,21,24,27,30]\times 10^{3}$ trajectories in order to study the relationship between the size of the batch dataset and the quality of the policies learned by FQI-log and FQI-sq. We move the successful trajectories to the front of the batch dataset so that they are included for all values of $n$.
@@
Since BFGS does not require tuning and the same batch datasets are given to FQI-log and FQI-sq, the only variable effecting the performance of the two methods is the loss being minimized. As shown in Fig. 1, FQI-log accumulates much smaller cost then FQI-sq using fewer samples, irrespective of the number of trajectories that reach the top of the hill. FQI-log is also able to learn a near-optimal policy with only a single successful trajectory. In batch RL collecting good trajectories is often expensive, so making efficient use of the few that appear in the batch dataset is an attractive algorithmic feature. As the number of trajectories that reach the top of the hill increases, so too does the performance of FQI-sq. Since the optimal value on this problem is zero, FQI-log is able to learn a near-optimal policy using fewer samples than FQI-sq.

### v5

@@ 4 FQI-log: Fitted Q-Iteration with log-loss
Input: A dataset $D_{n}=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i\in[n]}$, a function class $\mathcal{F}\subseteq[0,1]^{\mathcal{S}\times\mathcal{A}}$ and $k\in\mathbb{N}$.
@@
    $f_{j}\leftarrow\mathop{\rm arg\,min}\limits_{f\in\mathcal{F}}\sum_{i=1}^{n}\ell_{\log}\left(f(S_{i},A_{i}),C_{i}+\gamma f_{j-1}^{\wedge}(S_{i}^{\prime})\right)$
@@ 5 Theoretical Results
In this section, we present our main theoretical contribution, the first small-cost bound for an efficient algorithm in batch RL.
@@
size=,color=orange!20!white,]SR: will we use $D$ or $D_{n}$? Both are used in different places and I have no strong preference. CS; $D_{n}$. Given a dataset $D_{n}=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ with $n\in\mathbb{N}$ and a finite function class $\mathcal{F}\subseteq[0,1]^{\mathcal{S}\times\mathcal{A}}$ that satisfy Assumptions 3.1, 3.2, 3.3 and 3.4, it holds with probability $1-\delta$ that the suboptimality gap $g=\bar{v}^{\pi_{k}}-\bar{v}^{\star}$ of the output policy of FQI-log after $k$ iterations, $\pi_{k}=\pi_{f_{k}}$, satisfies
$g$ $\leq\tilde{\mathcal{O}}\left(\frac{1}{(1-\gamma)^{2}}\left(\sqrt{\frac{\bar{v}^{\star}CN}{n}}+\frac{CN}{(1-\gamma)^{2}n}+\gamma^{k}\right)\right)\,,$
where $N=\log(|\mathcal{F}|/\delta)$ and $C$ is defined in Assumption 3.2. size=,color=red!20!white,]Cs: missing function set complexity terms
The full statement of Theorem 5.1, including lower order terms, can be found in Appendix B along with its proof. Compared to prior error bounds for FQI (Antos et al., 2007, 2008; Munos & Szepesvári, 2008; Farahmand, 2011; Lazaric et al., 2012; Chen & Jiang, 2019), to the best of our knowledge, Theorem 5.1 is the first that contains the instance-dependent optimal cost $\bar{v}^{\star}$. This makes Theorem 5.1 a small-cost bound, also referred to as a first-order (Freund & Schapire, 1997; Neu, 2015) or small-loss (Lykouris et al., 2022; Wang et al., 2023) bound in the learning theory literature. All previous results for FQI obtain an error bound independent of $\bar{v}^{\star}$, and cannot be made to scale with $\bar{v}^{\star}$ due to their use of squared loss (see Theorem 2 of Foster & Krishnamurthy (2021), mentioned earlier). Finally, we highlight that Theorem 5.1 is the first small-cost bound for a batch RL algorithm that is computationally efficient when efficient regression oracles are available, as is the case when $\mathcal{F}$ is the set of logit models with weights bounded in $2$-norm. While technically this is outside of the scope of Theorem 5.1 (since in its current form this result covers only finite model classes), with some extra work and with appropriate modifications one can show that Theorem 5.1 continues to hold for infinite model classes, such as the mentioned logit class.
@@ 5.1 Proof Sketch
The purpose of this section is to give a sketch of the proof of Theorem 5.1. For the full proof, see Appendix B. We start by defining the pointwise triangular deviation of $f$ from $q^{\star}$,
@@
which is closely related to triangular discrimination (Topsøe, 2000). We can relate $\Delta^{2}_{f}$ to the Hellinger distance via the following lemma:
@@
where for $p=q=0$ we define the left-hand side to be zero and $\operatorname{h}^{2}(p\mathrel{\|}q)=\frac{1}{2}(\sqrt{p}-\sqrt{q})^{2}+\frac{1}{2}(\sqrt{1-p}-\sqrt{1-q})^{2}$ is the squared Hellinger distance. size=,color=red!20!white,]Cs: big bug was here.. watch out guys!
The proof of Lemma 5.2 is deferred to Section A.1. The idea to relate the pointwise triangular deviation to the squared Hellinger distance was first employed by Foster & Krishnamurthy (2021) in analyzing regret bounds for contextual bandits. Our proof can be summarized by the following three main steps, which correspond to the three terms given in Lemma 5.2.
@@
The first step in the proof is to decompose the error (or suboptimality gap), $\bar{v}^{\pi_{k}}-\bar{v}^{\star}$, into the product of a small-cost term and the pointwise triangular deviation of $f$ from $q^{\star}$. The analysis in this step is inspired by the proof of Lemma 1 of Foster & Krishnamurthy (2021). We deviate from their analysis to avoid introducing an extra $|\mathcal{A}|$ factor in the bound. We use the performance difference lemma (Lemma B.4), a multiplicative Cauchy-Schwarz (Lemma B.5), i.e. for distribution $\nu$
@@
and an implicit inequality (i.e. Lemma B.7, step $\star$ in the proof of Proposition B.2) to get a small-cost decomposition of the error:
@@
$\bar{v}^{\pi}-\bar{v}^{\star}\leq\tilde{C}\left(\frac{D_{f}}{1-\gamma}\sqrt{\bar{v}^{\star}}+\frac{D_{f}^{2}}{(1-\gamma)^{2}}\right)\,.$
where $\tilde{C}>0$ is an absolute constant.
@@
where $\nu\in\mathcal{M}_{1}(\mathcal{S}\times\mathcal{A})$. In Lemma B.13 we next establish that $\mathcal{T}$ is a $\gamma$-pseudo-contraction at $q^{\star}$ with respect to Hellinger distances: For any $f\geq 0$, $\nu\in\mathcal{M}_{1}(\mathcal{S}\times\mathcal{A})$,
$\left\|\sqrt{\mathcal{T}f}-\sqrt{\mathcal{T}q^{\star}}\right\|_{2,\nu}\leq\gamma^{1/2}\left\|\sqrt{f}-\sqrt{q^{\star}}\right\|_{2,\nu^{\prime}}\,,$
@@
Contraction arguments have a long history (Bertsekas, 1995; Littman, 1996; Antos et al., 2007; Chen & Jiang, 2019) in the analysis of dynamic programming algorithms that solve MDPs. The novelty is that we needed to bound the pointwise triangular deviation, which led to new analysis with Hellinger distances.
In Lemma B.15, the combination of a standard change of measure argument (that uses the definition of the concentration coefficient $C$) and a standard contraction argument, we get
@@
where $\mu$ is the exploratory distribution in Assumption 3.1. In batch RL, we often want to learn $q^{\star}$, or a policy whose value is close to $q^{\star}$. However, when using function approximation FQI, the regressor corresponding to the training data is $\mathcal{T}f$. Therefore, as demonstrated in Equation 6, if we can show the contraction property then we can bound the error between $f$ and $q^{\star}$ by the error between $f$ and $\mathcal{T}f$. As we will argue shortly, the error between $f$ and $\mathcal{T}f$ goes to zero as the size of the batch dataset grows.
Step 3: Error control/propagation
@@
Then by application of Theorem A.3, we have that if $f$ is the minimizer of $\ell_{\log}$ with respect to the batch dataset $D_{n}$, then
$\sqrt{2}\left\|\operatorname{h}^{2}(f\mathrel{\|}\mathcal{T}f)\right\|_{1,\mu}^{1/2}\lesssim\frac{2}{\sqrt{n}}\,,$
where we use $\lesssim$ informally to highlight the most salient elements of the inequality. We then combine all the results to control the pointwise triangular deviation in Proposition 5.3, i.e $D_{f}$,
@@
This provides a sketch for proving Theorem 5.1. In the full proof we also need to control each iterate of FQI-log, i.e. the $f_{j}$’s. We fill in the missing details in our appendix.
@@ 6 Numerical Experiments
The goal of our experiments is to provide insights into the benefits of using FQI-log for learning a near-optimal policy in batch reinforcement learning. We run our first set of experiments in reinforcement learning with logit models, as this setting allows us to best compare FQI-log to FQI-sq without other confounding factors. For these experiments, we used two standard control tasks; “mountain car” and “inverted pendulum”. The tasks are set up as fixed horizon episodic problems where at the end of an episode, if the goal is met no cost is incurred, otherwise a cost of one is incurred. The two environments differ in that in one of them the goal region is small, in the other the goal region is large. In both tasks, some policies can reach the goal, but many fail. As it is known that the goals in these problems can be met, we expect FQI-log to do better than FQI-sq on these environments.
Our second set of experiments aim at verifying whether the recommendation to switch to log loss transfers to deep RL (DRL), i.e., to more complex function classes, regression methods and environments. For these experiments, we started from the work of Agarwal et al. (2020), who tested various DRL methods, including C51 of (Bellemare et al., 2017), a distributional RL algorithm, which was found to be one of the most capable of the methods tested. As noted in the introduction, Wang et al. (2023) showed a small-cost bound for a distributional RL method; hence, our research question is whether a simpler log-loss based method can compete with these (more complex) distributional RL algorithms. To create a real challenge, we picked the two environments (Asterix and Sequest) from Agarwal et al. (2020) where C51 significantly outperformed DQN-sq, the DRL version of FQI and copied their setting.
6.1 Aiming for a Goal: Mountain Car
@@
Figure 1: The value of the policy learned by FQI as a function of the size of batch dataset. The results are averaged over $90$ independently collected datasets. The figures on the left, middle and right are generated using batch data that contain only $1$, $5$, and $30$ successful trajectories respectively. The standard error of the mean is reported via the shaded region.
We first evaluate FQI-log and FQI-sq on an episodic sparse cost variant of mountain car with episodes lasting for $800$ steps. (While we showed our results for the discounted setting, they are expected to hold in episodic problems as well, with small modifications.) Following Moore (1990), this environment consists of a $2$-dimensional continuous state space of $[-1.2,0.6]\times[-0.07,0.07]$ and $3$ discrete actions; the states represent a position and velocity of an underpowered car that can be accelerated left, right, or not accelerated, until the top of a hill is reached when the dynamics is turned on, and the car remains in place regardless of the actions. The cost is 0 at all timesteps except the last, when a cost of 1 is received if the learner has not reached the hilltop. We consider the undiscounted version of the problem (i.e., $\gamma=1$). An optimal policy for this setting reaches zero cost if it reliably reaches the top of the hill in $800$ steps or less, regardless of the exact time. For $\eta_{1}$, the initial state distribution, we use a Dirac that outs the car at the bottom of the hill with zero velocity with probability one.
@@
The feature vectors assigned to states are $16$ dimensional and come from a Fourier basis of order $4$, following Konidaris et al. (2011) and Chapter 9 of Sutton & Barto (2018). With this, for time step $h\in[800]$, the estimator uses $\theta_{h}=(\theta_{a}^{h})_{a\in\mathcal{A}}\in\mathbb{R}^{48}$ to produce the estimate $f_{h}(s,a)=\sigma(\left<\phi(s),\theta_{a}^{h},)\right>$, where $\sigma(x)=(1+\exp(-x))^{-1}$ is the sigmoid function. This variant of FQI-log with sigmoid functions is closely related to the logistic temporal-difference learning algorithm proposed in Appendix A of the PhD thesis of Silver (2009). We employ the BFGS method, a quasi-Newton method with no learning rate, to find the minimizer of the losses. While for strongly-convex functions, BFGS is known to converge to the global minimum superlinearly (Dennis & Moré, 1974), there is no guarantee that BFGS will find a global minimum when the loss is the squared loss. Finally, each batch dataset is constructed from a set of trajectories collected by running the uniform random policy from the initial state $30,000$ times. We use rejection sampling in order to guarantee each dataset has $i$ trajectories that reach the top of the hill with $i\in\{1,5,30\}$. We train both FQI-log and FQI-sq on the same batch datasets with the first $n=[1,3,6,9,12,15,18,21,24,27,30]\times 10^{3}$ trajectories in order to study the relationship between the size of the batch dataset and the quality of the policies learned by FQI-log and FQI-sq. We move the successful trajectories to the front of the batch dataset so that they are included for all values of $n$.
@@
Since BFGS does not require tuning and the same batch datasets are given to FQI-log and FQI-sq, the only variable effecting the performance of the two methods is the loss being minimized. As shown in Figure 1, FQI-log accumulates much smaller cost then FQI-sq using fewer samples, irrespective of the number of trajectories that reach the top of the hill. FQI-log is also able to learn a near-optimal policy with only a single successful trajectory. In batch RL collecting good trajectories is often expensive, so making efficient use of the few that appear in the batch dataset is an attractive algorithmic feature. As the number of trajectories that reach the top of the hill increases, so too does the performance of FQI-sq. Since the optimal value on this problem is zero, FQI-log is able to learn a near-optimal policy using fewer samples than FQI-sq.
6.2 Avoiding Failure: Inverted Pendulum
We further evaluate FQI-log and FQI-sq on the inverted pendulum environment (Lagoudakis & Parr, 2003; Riedmiller, 2005), where the goal is to balance an inverted pendulum, by applying forces to it. The state space is two dimensional (angle, angular velocity) and there are three actions (push left, right, no change). The environment dynamics are as described by Lagoudakis & Parr (2003), except that (i) when the pendulum falls below horizontal, the state is frozen there and (ii) we clip the angular momentum to be in $[-5,5]$ instead of letting it take arbitrary real values; the clipping is done to facilitate the use of Fourier basis features, which we found works better than the radial basis functions used by Lagoudakis & Parr (2003). As we were using order $4$ Fourier features, the parameter space in this case is $4\times 4\times 3=48$ dimensional. The same logit class was used as in the case of mountain car; for fitting BFGS was used. The cost structure is as follows: The cost of letting the pendulum fall bellow the horizontal is 00. There is no cost otherwise. Again, we know there exist policies that achieves near-zero cost. The discount factor is $\gamma=0.95$. All datasets are collected by a policy that selects actions uniformly at random until failure (which happens typically after $6$ steps) starting from a close to upright, random position. We performed $90$ independent trials, each trial consisting of fitting FQI-log and FQI-sq on the same data for $k=300$ rounds. We then evaluate the learned policies $1000$ times on whether the inverted pendulum is still balanced after $3000$ steps, starting from a near upright, random position. Results are shown in Figure 2. As with the mountain car experiments, we see that FQI-log uses fewer samples than FQI-sq to learn a good policy.
Figure 2: The portion of the time that a policy learned by FQI was able to balance the pendulum for $3000$ steps. Results are averaged over $90$ independently collected datasets and each learned policy is tested on $1000$ initializations. The standard error of the mean is shaded.